Let R = Z + √2 Z and S = the ring of 2 x 2
matrices of the form [
a 2b
b. a
. ] Where a,b belongs to Z

Show that f : R --> S defined by
f (a + √2 b) = [
a. 2b
b. a
] is an isomorphism of rings.
1
Expert's answer
2020-11-24T16:53:11-0500

Surjectivity is obvious since S is the set of matrices of that special form.

Injectivity: "f(a+\sqrt{2}b)=f(c+\sqrt{2}d)". Hence, equating the columns of the matrices we get, "a=c,b=d." Hence injective.

Homomorphism: "f[(a+\sqrt{2}b)+(c+\sqrt{2}d)]=f[a+c+\sqrt{2}(b+d)]""=\begin{bmatrix}
 a+c & 2b+2d \\
 b+d & a+c
\end{bmatrix}." "=\begin{bmatrix}
 a & 2b \\
 b & a
\end{bmatrix} +\begin{bmatrix}
 c & 2d \\
 d & c
\end{bmatrix} = f(a+\sqrt{2}b)+f(c+\sqrt{2}d)." Also "f[(a+\sqrt{2}b)(c+\sqrt{2}d)]=f[ac+2bd+\sqrt{2}(bc+ad)]" "=\begin{bmatrix}
 ac+2bd & 2bc+2ad \\
 bc+ad & ac+2bd
\end{bmatrix}." "=\begin{bmatrix}
 a & 2b \\
 b & a
\end{bmatrix}\begin{bmatrix}
 c & 2d \\
 d & c
\end{bmatrix}=" "f(a+\sqrt{2}b)f(c+\sqrt{2}d)." Also "f(1)=\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}" . Hence homomorphism